Write the equation of the line that is parallel to the line 2x-3y= 0 and contains the point (6,1)
Answer by nerdybill(7384)

Write the equation of the line that is parallel to the line 2x-3y= 0 and contains the point (6,1)
.
First, determine the slop of:
2x-3y= 0
by putting it into the "slope-intercept" form
-3y= -2x
y= (2/3)x
.
So, slope is 2/3. A line that is parallel must have the same slope.
.
Plug the slope (2/3) and the given point(6,1) into the "point-slope" form:
y - y1 = m(x - x1)
y - 1 = (2/3)(x - 6)
y - 1 = (2/3)x - 4
y = (2/3)x - 3 (this is what they're looking for)